# ClipTTA: Robust Contrastive Vision-Language Test-Time Adaptation

### v1

@@ 4.2 Model analysis
Table 4 presents an ablation of ClipTTA ’s components on four closed-set benchmarks. The first key observation is that the soft contrastive loss ($\mathcal{L}_{\text{s-cont}}$) alone accounts for the vast majority of the overall performance gains, demonstrating the importance of aligning the adaptation objective with CLIP’s pre-training. In low-accuracy settings, $\mathcal{L}_{\text{s-cont}}$ significantly outperforms TENT, achieving gains of +19.4 points on CIFAR-100-C and +22.7 points on ImageNet-C, where TENT even degrades CLIP’s performance. This confirms the vulnerability of entropy-based methods to pseudo-label drift and class collapse and supports the enhanced robustness of $\mathcal{L}_{\text{s-cont}}$, as theoretically analyzed in Sec. 3.2. On average, $\mathcal{L}_{\text{s-cont}}$ improves over TENT by +11.4 points. Adding the regularization loss ($\mathcal{L}_{\text{reg}}$) further improves overall results by +0.3 pts, while incorporating the confident memory ($\mathcal{M}$) brings additional +0.8 pts gains.
@@
Figure 4 provides empirical insights supporting the gradient analysis in Sec. 3.2, by illustrating CLIPTTA’s stability and robustness over batches on CIFAR-10-C. ClipTTA is the only method that steadily improves accuracy throughout adaptation while all competing objectives plateau or degrade (Fig.4a). This stability is closely linked to CLIPTTA’s ability to maintain prediction diversity. As shown in Fig.4b, prediction entropy remains nearly constant for ClipTTA, whereas TENT exhibits a sharp drop in entropy, indicating collapse toward a small subset of classes. This behavior results in harmful label drift. Fig. 4c tracks the deterioration ratio, defined as the fraction of initially correct predictions that become incorrect during adaptation. TENT reaches over 30% deterioration, compared to less than 7% with ClipTTA. These findings confirm the significant stabilizing effect of our batch-aware contrastive loss during adaptation.
@@ A.1 Gradient Analysis
[…]
@@
with $q_{ik}$ the probability of image ${\bm{x}}_{i}$ being classified as class $k$ (see Eq. 1). The gradient of $\mathcal{L}_{\text{TENT}}$ w.r.t. ${\bm{z}}_{i}$ is:
$\begin{split}\nabla_{{\bm{z}}_{i}}\mathcal{L}_{\text{TENT}}=-\sum_{k=1}^{C}\nabla_{{\bm{z}}_{i}}q_{ik}\log{q_{ik}}&=-\sum_{k=1}^{C}(1+\log{q_{ik}})\nabla_{z_{i}}q_{ik}\\ &=-\sum_{k=1}^{C}(1+\log{q_{ik}})~{}q_{ik}\sum_{c=1}^{C}q_{ic}~{}({\bm{z}}_{t}^{k}-{\bm{z}}_{t}^{c})\\ &=-\sum_{k=1}^{C}~{}\big{[}\sum_{c=1}^{C}\log{\frac{q_{ik}}{q_{ic}}}~{}q_{ic}~{}\big{]}~{}q_{ik}~{}{\bm{z}}_{t}^{k}\\ \end{split}$ (10)
From Eq. 10 we can see that the gradient will always push ${\bm{z}}_{i}$ in the direction of the predicted class $\hat{k}$ because in that case we have $\log{\frac{q_{i\hat{k}}}{q_{ic}}}>0,\forall c\neq\hat{k}$. And there is no mechanism allowing to reduce the magnitude of the gradient towards the predicted class even when we are approaching a situation of collapse.
@@
From Eq. 11, we observe that CLIP’s contrastive loss consistently drives the visual embedding ${\bm{z}}_{i}$ toward the embedding of its predicted class $\widehat{{\bm{z}}_{t}}^{i}$, as $w_{k,i}\leq 1$. However, the gradient’s magnitude is influenced by the proportion of predictions assigned to the same class within the batch. Specifically, as the system approaches a collapse scenario (i.e., $w_{k,i}\rightarrow 1$), the gradient of $\mathcal{L}_{\text{cont.}}$ diminishes and eventually vanishes:
$||\nabla_{z_{i}}\mathcal{L}_{\text{cont.}}||\underset{w_{k,i}\rightarrow 1}{\rightarrow}0$ (12)
@@
$\mathcal{L}_{\text{s-cont}}=$ $\sum_{i=1}^{N}-\sum_{j=1}^{N}p(\hat{{\bm{t}}}_{j}|{\bm{x}}_{i})\log{p(\hat{{\bm{t}}}_{j}|{\bm{x}}_{i})}$
@@
Now we can use the fact that $\nabla_{z_{i}}-\log{p(\hat{{\bm{t}}}_{j}|{\bm{x}}_{i})}=-\widehat{{\bm{z}}_{t}}^{j}+\sum_{k=1}^{C}w_{k,i}~{}{\bm{z}}_{t}^{k}$ based on the computation $\mathcal{L}_{\text{cont.}}$ in Eq. 11. Therefore, we have:
@@
The gradient of $\mathcal{L}_{\text{s-cont}}$ does not solely push the visual embedding toward the predicted class. Instead, it incorporates other predictions within the batch to guide the gradient direction, thereby mitigating the risk of pseudo-labeling errors. However, similar to CLIP’s contrastive loss, the gradient diminishes as we approach a collapse scenario. In the case of collapse, where all examples in the batch are predicted to belong to the same class $c$ , the following conditions hold: $w_{c}({\bm{x}}_{i})=1$ and $w_{k,i}=0,\forall k\neq c$, and $\widehat{{\bm{z}}_{t}}^{j}={\bm{z}}_{t}^{c}\forall j$. Consequently, the term $[-\widehat{{\bm{z}}_{t}}^{j}+\sum_{k=1}^{C}w_{k,i}~{}{\bm{z}}_{t}^{k}]$ cancels out, leading to a null gradient.
@@
We derive Eq. 7 in the main paper, starting from Eq. 13, and assuming that the classification task comprises two classes $K=\{a,b\}$, with $N=N_{a}+N_{b}$ as the total batch size. To build on the intuition of the working mechanisms of our soft contrastive loss, we adopt the case where class $a$ is dominant in the batch (i.e., $N_{a}\gg N_{b}$). First, we expand on the second sum term inside Eq. 13, as follows:
@@
We notice that we can partition the main sum term in Eq. 13 into two sums that account for the $N_{a}$ samples predicted as class $a$, and the $N_{b}$ samples predicted as class $b$:
@@
From Eq. 17, we observe that the gradient is influenced by the ratios $\log{\frac{\bar{q}_{k}}{\bar{q}_{j}}}$, driving it towards the classes that are underrepresented in the batch predictions. The use of the regularization loss in conjunction with our soft contrastive loss creates a powerful combined effect, enabling the effective relabeling of misclassified examples, as discussed in Appendix B.
[…]
@@
[…]
@@ A.2 Analysis of OCE Loss
As discussed in the main paper, our outlier contrastive exposure (OCE) in Eq. 9 of the main paper is a special case of the intra-class variance minimization:
@@
with $p_{\text{id}}=\frac{1}{N}\sum_{i}w_{i}$ and $p_{\text{ood}}=\frac{1}{N}\sum_{i}(1-w_{i})$. This is further condensed into the loss function Eq. 21:
@@ Appendix B Discussion on ClipTTA’s robustness
Mitigating pseudo-label errors.  In Fig. 5, we present a controlled toy example demonstrating how ClipTTA effectively mitigates misclassifications. This example features a batch of six samples in a three-class classification problem. It focuses on the gradient orientations of the TENT, ClipTTA, and regularized ClipTTA losses for a single misclassified and ambiguous sample. The sample in question exhibits high probabilities for both the predicted and correct labels, indicating low confidence. The gradient of the ClipTTA loss is directed toward the correct label, working to minimize the difference between the top two probabilities—a behavior further amplified by the regularized ClipTTA loss. In contrast, TENT prioritizes increasing the highest probability, thereby reinforcing the incorrect prediction.
@@
We provide quantitative insights on the CIFAR-10-C dataset in Fig. 6. In Fig. 6-a, we observe the collapse of TENT, while ClipTTA maintains robust performance. To further analyze this, we quantify two key metrics: the “improvement ratio” shown in Fig. 6-b, which represents the proportion of misclassified examples that are correctly classified after adaptation, and the “deterioration ratio” shown in Fig. 6-c, which denotes the proportion of correctly classified examples that become misclassified after adaptation. ClipTTA outperforms TENT by achieving a higher improvement ratio and a lower deterioration ratio.
@@ C.2 Details on baselines and datasets
We provide more detailed information about the datasets that compose the different benchmarks used through the main paper. For all the experiments, images of different sizes were reshaped for compatibility with CLIP (i.e., to 224$\times$224).
@@
We employed CIFAR-10 and CIFAR-100 [31], both containing 10,000 images of size 28$\times$28, and spanning 10 and 100 classes, respectively. We use Imagenet [32] as a larger-scale dataset, with 1000 classes and 50,000 images in total.
@@ C.3 Extended experimental results
We conduct additional runs of ClipTTA to assess its sensitivity to random initialization, reporting the mean accuracy and 95% confidence interval in Tab. 16. The results indicate that ClipTTA is highly stable, with very low variance across independent runs. The tight confidence intervals (e.g., ±$\pm 0.01$ on ImageNet) confirm the reliability and reproducibility of the observed performance gains, further supporting the robustness of the method across different datasets.

### v2

@@ 4.2 Model analysis
Table 4 presents an ablation of ClipTTA ’s components on four closed-set benchmarks. The first key observation is that the soft contrastive loss ($\mathcal{L}_{\text{s-cont}}$) alone accounts for the vast majority of the overall performance gains, demonstrating the importance of aligning the adaptation objective with CLIP’s pre-training. In low-accuracy settings, $\mathcal{L}_{\text{s-cont}}$ significantly outperforms TENT, achieving gains of +19.4 points on CIFAR-100-C and +22.7 points on ImageNet-C, where TENT even degrades CLIP’s performance. This confirms the vulnerability of entropy-based methods to pseudo-label drift and class collapse and supports the enhanced robustness of $\mathcal{L}_{\text{s-cont}}$, as theoretically analyzed in Sec.˜3.2. On average, $\mathcal{L}_{\text{s-cont}}$ improves over TENT by +11.4 points. Adding the regularization loss ($\mathcal{L}_{\text{reg}}$) further improves overall results by +0.3 pts, while incorporating the confident memory ($\mathcal{M}$) brings additional +0.8 pts gains.
@@
Figure 4 provides empirical insights supporting the gradient analysis in Sec.˜3.2, by illustrating CLIPTTA’s stability and robustness over batches on CIFAR-10-C. ClipTTA is the only method that steadily improves accuracy throughout adaptation while all competing objectives plateau or degrade (Fig.4a). This stability is closely linked to CLIPTTA’s ability to maintain prediction diversity. As shown in Fig.4b, prediction entropy remains nearly constant for ClipTTA, whereas TENT exhibits a sharp drop in entropy, indicating collapse toward a small subset of classes. This behavior results in harmful label drift. Fig.˜4c tracks the deterioration ratio, defined as the fraction of initially correct predictions that become incorrect during adaptation. TENT reaches over 30% deterioration, compared to less than 7% with ClipTTA. These findings confirm the significant stabilizing effect of our batch-aware contrastive loss during adaptation.
@@ A.1 Gradient Analysis
In this section, we provide a detailed analysis of the gradients of the TENT loss $\mathcal{L}_{\text{TENT}}$, CLIP’s contrastive loss $\mathcal{L}_{\text{cont.}}$ (i.e. using hard pseudo-captions), our soft contrastive loss $\mathcal{L}_{\text{s-cont.}}$, the regularization loss $\mathcal{L}_{\text{reg}}$ and the ClipTTA loss $\mathcal{L}_{\text{C{lip}TTA}}$. Furthermore, we show how, benefiting from the information of other predictions in the batch, both contrastive losses allow to avoid collapse. Finally, when combined with the regularization loss, ClipTTA allows mitigating the effect of pseudo label errors.
@@
with $q_{ik}$ the probability of image ${\bm{x}}_{i}$ being classified as class $k$ (see Eq.˜1). The gradient of $\mathcal{L}_{\text{TENT}}$ w.r.t. ${\bm{z}}_{i}$ is:
[…]
From Eq.˜10 we can see that the gradient will always push ${\bm{z}}_{i}$ in the direction of the predicted class $\hat{k}$ because in that case we have $\log{\frac{q_{i\hat{k}}}{q_{ic}}}>0,\forall c\neq\hat{k}$. And there is no mechanism allowing to reduce the magnitude of the gradient towards the predicted class even when we are approaching a situation of collapse.
@@
From Eq.˜11, we observe that CLIP’s contrastive loss consistently drives the visual embedding ${\bm{z}}_{i}$ toward the embedding of its predicted class $\widehat{{\bm{z}}_{t}}^{i}$, as $w_{k,i}\leq 1$. However, the gradient’s magnitude is influenced by the proportion of predictions assigned to the same class within the batch. Specifically, as the system approaches a collapse scenario (i.e., $w_{k,i}\rightarrow 1$), the gradient of $\mathcal{L}_{\text{cont.}}$ diminishes and eventually vanishes:
[…]
@@
$\mathcal{L}_{\text{s-cont}}=$ $\sum_{i=1}^{N}-\sum_{j=1}^{N}p(\hat{{\bm{t}}}_{j}|{\bm{x}}_{i})\log{p(\hat{{\bm{t}}}_{j}|{\bm{x}}_{i})}\$
@@
Now we can use the fact that $\nabla_{z_{i}}-\log{p(\hat{{\bm{t}}}_{j}|{\bm{x}}_{i})}=-\widehat{{\bm{z}}_{t}}^{j}+\sum_{k=1}^{C}w_{k,i}~{\bm{z}}_{t}^{k}$ based on the computation $\mathcal{L}_{\text{cont.}}$ in Eq.˜11. Therefore, we have:
@@
The gradient of $\mathcal{L}_{\text{s-cont}}$ does not solely push the visual embedding toward the predicted class. Instead, it incorporates other predictions within the batch to guide the gradient direction, thereby mitigating the risk of pseudo-labeling errors. However, similar to CLIP’s contrastive loss, the gradient diminishes as we approach a collapse scenario. In the case of collapse, where all examples in the batch are predicted to belong to the same class $c$ , the following conditions hold: $w_{c}({\bm{x}}_{i})=1$ and $w_{k,i}=0,\forall k\neq c$, and $\widehat{{\bm{z}}_{t}}^{j}={\bm{z}}_{t}^{c}\forall j$. Consequently, the term $\widehat{{\bm{z}}_{t}}^{j}+\sum_{k=1}^{C}w_{k,i}~{\bm{z}}_{t}^{k}]$ cancels out, leading to a null gradient.
@@
We derive Eq.˜7 in the main paper, starting from Eq.˜13, and assuming that the classification task comprises two classes $K=\{a,b\}$, with $N=N_{a}+N_{b}$ as the total batch size. To build on the intuition of the working mechanisms of our soft contrastive loss, we adopt the case where class $a$ is dominant in the batch (i.e., $N_{a}\gg N_{b}$). First, we expand on the second sum term inside Eq.˜13, as follows:
@@
We notice that we can partition the main sum term in Eq.˜13 into two sums that account for the $N_{a}$ samples predicted as class $a$, and the $N_{b}$ samples predicted as class $b$:
@@
From Eq.˜17, we observe that the gradient is influenced by the ratios $\log{\frac{\bar{q}_{k}}{\bar{q}_{j}}}$, driving it towards the classes that are underrepresented in the batch predictions. The use of the regularization loss in conjunction with our soft contrastive loss creates a powerful combined effect, enabling the effective relabeling of misclassified examples, as discussed in Appendix˜B.
Gradient of $\mathcal{L}_{\text{C{lip}TTA}}$.
@@
$\mathcal{L}_{\text{C{lip}TTA}}=\mathcal{L}_{\text{s-cont}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (18)
Therefore the gradient of $\mathcal{L}_{\text{C{lip}TTA}}$ writes:
$\begin{split}\nabla_{z_{i}}\mathcal{L}_{\text{C{lip}TTA}}&=\sum_{j=1}^{N}\beta_{i,j}[-\widehat{{\bm{z}}_{t}}^{j}+\sum_{k=1}^{C}w_{k,i}~{\bm{z}}_{t}^{k}]+\lambda_{reg}\frac{1}{N}\sum_{k=1}^{C}[\sum_{j=1}^{C}q_{ij}\log{\frac{\bar{q}_{k}}{\bar{q}_{j}}}]q_{ik}~{\bm{z}}_{t}^{k}.\end{split}$ (19)
Depending on the composition of the batch, we can see that $\mathcal{L}_{\text{C{lip}TTA}}$ will strongly benefit from the contribution of the soft-contrastive loss to provide accurate adaptation, or will be able to correct misclassified examples due to the positive interaction of the combined corrective terms in $\nabla_{z_{i}}\mathcal{L}_{\text{s-cont}}$ and $\nabla_{z_{i}}\mathcal{L}_{reg}$.
@@ A.2 Analysis of OCE Loss
As discussed in the main paper, our outlier contrastive exposure (OCE) in Eq.˜9 of the main paper is a special case of the intra-class variance minimization:
@@
with $p_{\text{id}}=\frac{1}{N}\sum_{i}w_{i}$ and $p_{\text{ood}}=\frac{1}{N}\sum_{i}(1-w_{i})$. This is further condensed into the loss function Eq.˜21:
@@ Appendix B Discussion on ClipTTA’s robustness
Mitigating pseudo-label errors.  In Fig.˜5, we present a controlled toy example demonstrating how ClipTTA effectively mitigates misclassifications. This example features a batch of six samples in a three-class classification problem. It focuses on the gradient orientations of the TENT, ClipTTA, and regularized ClipTTA losses for a single misclassified and ambiguous sample. The sample in question exhibits high probabilities for both the predicted and correct labels, indicating low confidence. The gradient of the ClipTTA loss is directed toward the correct label, working to minimize the difference between the top two probabilities—a behavior further amplified by the regularized ClipTTA loss. In contrast, TENT prioritizes increasing the highest probability, thereby reinforcing the incorrect prediction.
@@
We provide quantitative insights on the CIFAR-10-C dataset in Fig.˜6. In Fig.˜6-a, we observe the collapse of TENT, while ClipTTA maintains robust performance. To further analyze this, we quantify two key metrics: the “improvement ratio” shown in Fig.˜6-b, which represents the proportion of misclassified examples that are correctly classified after adaptation, and the “deterioration ratio” shown in Fig.˜6-c, which denotes the proportion of correctly classified examples that become misclassified after adaptation. ClipTTA outperforms TENT by achieving a higher improvement ratio and a lower deterioration ratio.
@@ C.2 Details on baselines and datasets
We provide more detailed information about the datasets that compose the different benchmarks used through the main paper. For all the experiments, images of different sizes were reshaped for compatibility with CLIP (i.e., to $\times224$).
@@
We employed CIFAR-10 and CIFAR-100 [31], both containing 10,000 images of size $\times28$, and spanning 10 and 100 classes, respectively. We use Imagenet [32] as a larger-scale dataset, with 1000 classes and 50,000 images in total.
@@ C.3 Extended experimental results
We conduct additional runs of ClipTTA to assess its sensitivity to random initialization, reporting the mean accuracy and 95% confidence interval in Tab.˜16. The results indicate that ClipTTA is highly stable, with very low variance across independent runs. The tight confidence intervals (e.g., $\pm 0.01$ on ImageNet) confirm the reliability and reproducibility of the observed performance gains, further supporting the robustness of the method across different datasets.